OpenStudy (anonymous):

will fan and medal.. I REALLY NEED HELP The sum of the measures of angle X and Y is 90. If the measures of angle X is 30 less than twice the measure of angle Y, what is the measure of angle X? A. 20 B. 35 C. 50 D.65

Directrix (directrix):

Let x and y be the measures of angles X and Y. The sum of the measures of angle X and Y is 90. So, x = y = 90 angle X is 30 less than twice the measure of angle Y, So, x = 2y -30

OpenStudy (anonymous):

x + y = 90 x = 2y -30 substitute x in the first equation (2y-30) + y = 90 solve for y, then plug in y in the first equation to find x

OpenStudy (anonymous):

i think its x= 30-2y actually.

OpenStudy (anonymous):

so then how do i get x from there?

OpenStudy (anonymous):

well they give you 2 equations to work with. one is x + y= 90 one is x = 30-2y

Directrix (directrix):

x + y = 90 ---> x = 90 - y x = 2y -30---->x = 2y-30 -----------

OpenStudy (anonymous):

plug in the 2nd equation into the first equation, then solve for y, after you solve for y you can solve for x by doing 90-y = x

OpenStudy (anonymous):

.....so what is x?

Directrix (directrix):

90-y = 2y - 30 y = 40 Subtract 40 from 90 to get x.

OpenStudy (anonymous):

you can't expect us to do your hw

OpenStudy (anonymous):

40-90=-50

OpenStudy (anonymous):

is that x?-50?

Directrix (directrix):

90-40 not 40 - 90

OpenStudy (anonymous):

so just 50?

OpenStudy (anonymous):

Can you help with this one? the yearbook club is having a bake sale to raise money for the senior class. LArge cupcakes are sold for 1.25 and small cupcakes are .75. If 105 cupcakes were sold for a total amount of 109.75, how many large cupcakes did the yearbook club sell?

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

@yoonlim19

Directrix (directrix):

Let L and S be the number of large and small cupcakes sold, respectively. Then, L + S= 105

OpenStudy (anonymous):

okay so then how do i figure out how many large cupcakes were sold?

OpenStudy (anonymous):

this question confuses me so much..D:

Directrix (directrix):

1.25L + .75S = 109.75 L + S= 105 ----------------------

Directrix (directrix):

The task is to solve those two equations simultaneously.

OpenStudy (anonymous):

okay. i know that but how do i use that to find out how many large cupcakes were sold

Directrix (directrix):

You use Algebra. 1.25L + .75S = 109.75 ----> multiply each member by 100 L + S= 105

Directrix (directrix):

125L +75S =109.75 L + S= 105 ---------------

Directrix (directrix):

125L +75S =109.75 L + S= 105 --> L = 105 - S

OpenStudy (anonymous):

confusion continues..

OpenStudy (anonymous):

okayy... so how do i get the amount of cupcakes with that?

Directrix (directrix):

125L +75S =10975 L + S= 105 --> L = 105 - S Solve these two equations in two unknowns simultaneously. Use the substitution method.

OpenStudy (anonymous):

Can you do that and like explain what your doing to get the answer so i can see how to do it?

Directrix (directrix):

Did you follow to this point: 125L +75S =10975 L + S= 105 --> L = 105 - S

OpenStudy (anonymous):

yeah

Directrix (directrix):

125L +75S =10975 125(105 -S) = 10975

OpenStudy (anonymous):

okay...

Directrix (directrix):

You multiply this --> 125(105 -S) and post what you get, okay?

OpenStudy (anonymous):

what is s?

OpenStudy (anonymous):

.75?

Directrix (directrix):

Just call it S

OpenStudy (anonymous):

13125s?

Directrix (directrix):

125(105 -S) = 13125 - 125S You have to use the Distributive Property

OpenStudy (anonymous):

okay..

OpenStudy (anonymous):

13000s?

Directrix (directrix):

125(105 -S) = 10975 13125 - 125S = 10975

OpenStudy (anonymous):

so whats L?

Directrix (directrix):

What is 10975 - 13125 = ?

OpenStudy (anonymous):

-2150 ^

Directrix (directrix):

- 125*S = -1250

Directrix (directrix):

What is -1250 divided by -125 =

OpenStudy (anonymous):

10

Directrix (directrix):

Something has run amok here because look at the Wolf's work: http://www.wolframalpha.com/input/?i=%28125L+%2B75S+%3D10975%29+and+%28L+%2B+S%3D+105%29

OpenStudy (anonymous):

that helps alot thanks

Directrix (directrix):

The Wolf has L = 62 and S = 43 so we have to back up and find where our error is.

Directrix (directrix):

The equations are correct but somewhere there is an error in our arithmetic.

OpenStudy (anonymous):

ok well ill look at that website.. thanks:D

Directrix (directrix):

Okay.